# Controllable Video Generation with Provable Disentanglement

### v2

@@ 1 Introduction
Video generation (vondrick2016generating, ; tulyakov2018mocogan, ; wang2022internvideo, ) has become a prominent research focus, driven by its wide-ranging applications in fields such as world simulators (openai2024video, ), autonomous driving (wen2024panacea, ; wang2023drivedreamer, ), and medical imaging (li2024endora, ; cao2024medical, ). In particular, controllable video generation (zhang2025moonshot, ) is essential for advancing more reliable and efficient video generation models. Despite the impressive results achieved by recent commercial or large-scale models such as Kling kling2024 and Wan wan2024wan , precise control over specific aspects of generated video remains a significant challenge, as illustrated in Figure 1. This issue may arise because these models typically represent the video as a unified 4D spatiotemporal block and apply conditioning signals (e.g., text prompts) directly to this global representation. Recent approaches (ho2022video, ; zhou2022magicvideo, ; yang2024cogvideox, ; opensora, ) follow a similar strategy, with differences in modeling frameworks (e.g., diffusion or VAE) and shapes (e.g., vectors or spatiotemporal blocks). However, these formulations often neglect the intricate spatio-temporal structure of videos, thereby limiting the ability to disentangle and control fine-grained factors, such as head movements and eye blinking in facial video generation.
@@
To address this issue, one intuitive solution is to learn a disentangled representation of the video, within which the internal relationships are often not considered. Some models (hyvarinen2000independent, ; tulyakov2018mocogan, ; yu2022generating, ; skorokhodov2022stylegan, ; wei2024dreamvideo, ) explicitly decompose video generation into two parts: motion and identity, representing dynamic and static information, respectively. This separation allows for more targeted control over each aspect, making it possible to modify the motion independently without affecting the identity. zhang2025moonshot ; shen2023mostganv leverage attention mechanisms to further disentangle different concepts within the video, enhancing the ability to control specific features with greater precision. fei2024dysen ; lin2023videodirectorgpt utilize Large Language Models to find the intricate temporal dynamics within the video and then enrich the scene with reasonable details, enabling a more transparent generative process. These methods are intuitive and effective, yet they lack a solid guarantee of disentanglement, making the control less predictable and potentially leading to unintentional coupling of different aspects of the video.
@@
These limitations of previous approaches motivate us to rethink the paradigm of video generation. Inspired by recent advancements in nonlinear Independent Component Analysis (ICA) (hyvarinen2017nonlinear, ; khemakhem2020variational, ; yao2022temporally, ; hyvarinen2016unsupervised, ) and the successful applications like video understanding (chen2024caring, ), we propose the Controllable Video Generative Adversarial Network (CoVoGAN) with a Temporal Transition Module plugin. Building upon StyleGAN2-ADA Karras2020ada , we distinguish between two types of factors: the dynamic factors that evolve over time, referred to as style dynamics, and the static factors that remain unchanged, which we call content elements. We also distinguish between different components within the dynamics, allowing for more precise control. By leveraging the minimal change principle, we demonstrate their block-wise identifiability (von2021self, ; li2024subspace, ) and find the conditions under which motion and identity can be disentangled, explaining the effectiveness of the previous line of video generation methods that separate motion and identity. In addition, we employ sufficient change property to disentangle different concepts of motion, such as head movement or eye blinking. Specifically, we introduce a flow (rezende2015variational, ) mechanism to ensure that the estimated style dynamics are mutually independent conditioned on the historical information. Furthermore, we prove the component-wise identifiability of the style dynamics and provide a disentanglement guarantee for the motion in the video.
@@
We conduct both quantitative and qualitative experiments on various video generation benchmarks. For quantitative evaluation, we use FVD (unterthiner2019fvd, ) to assess the quality of the generated videos. We also compare SAP kumar2017variational , modularity ridgeway2018learning and MCC to demonstrate the disentanglement capability. For qualitative analysis, we evaluate the degree of disentanglement by manipulating different dimensions of the latent variables across multiple datasets and comparing the resulting video outputs. Experimental results demonstrate that our method significantly outperforms other video generation models of a scale similar to CoVoGAN.
@@
  • We propose a Temporal Transition Module to achieve a disentangled representation, which leverages minimal change principle and sufficient change property.
@@ 2.1 Generating process
{$&{\mathbf{x}}_{t}=g({\mathbf{z}}_{t}^{s},{\mathbf{z}}^{c}),\\ &z_{t,i}^{s}=f^{s}_{i}\left(\mathbf{Pa}(z_{t,i}^{s}),\epsilon_{t,i}^{s}\right),\\ &z_{j}^{c}={f^{c}_{j}\left(\epsilon_{j}^{c}\right)},\end{aligned}\right.\text{with}\left\{\begin{aligned} \epsilon^{s}_{t,i}\sim p^{s}_{\epsilon_{i}},\\ \epsilon^{c}_{j}\sim p^{c}_{\epsilon_{j}},\end{aligned}\right.$ (1)
in which $z_{t,i}^{s},z_{i}^{c}\in\mathbb{R}$ refers to the $i$-th entry of ${\mathbf{z}}_{t}^{s}$ and $j$-th entry of ${\mathbf{z}}^{c}$, respectively. $\mathbf{Pa}(z_{t,i}^{s})$ refers to the time-deleyed parents of $z_{t,i}^{s}$. All noise terms $\epsilon_{t,i}$ and $\epsilon_{j}$ are independently sampled from their respective distributions: $p_{\epsilon_{i}}^{s}$ for the $i$-th entry of the style dynamics, and $p_{\epsilon_{j}}^{c}$ for the $j$-th entry of the content elements. The components of ${\mathbf{z}}_{t}^{s}$ are mutually independent, conditioned on all historical variables $\cup_{i=1}^{n_{s}}\mathbf{Pa}(z_{t,i}^{s})$. The non-parametric causal transition $f_{i}$ enables an arbitrarily nonlinear interaction between the noise term $\epsilon_{t,i}$ and the set of parent variables $\mathbf{Pa}(z_{t,i}^{s})$, allowing for flexible modeling of the style dynamics.
@@ 2.2 Identification of the latent causal process
$\ \ p_{(\hat{g},\hat{\textbf{f}}^{s},\hat{\textbf{f}}^{c},\hat{\textbf{p}}^{s},\hat{\textbf{p}}^{c})}(V)=p_{(g,\textbf{f}^{s},\textbf{f}^{c},\textbf{p}^{s},\textbf{p}^{c})}(V)$ (2)
$\Rightarrow$ $\ \ [{\mathbf{z}}_{t}^{s},{\mathbf{z}}^{c}]=h([\hat{{\mathbf{z}}}_{t}^{s},\hat{{\mathbf{z}}}^{c}]),{\mathbf{z}}^{c}=h_{c}(\hat{{\mathbf{z}}}^{c}),$
where []delimited-[][$\cdot$][ ⋅ ] denotes concatenation.
Illustration. When block-wise identification is achieved, content elements are effectively disentangled. As a result, motion control can be applied independently, allowing manipulation of motion without altering the video’s content. For example, in a video where a camera moves forward, it becomes easy to change the camera’s direction without affecting the scene itself.
@@
$\ \ p_{(\hat{g},\hat{\textbf{f}}^{s},\hat{\textbf{f}}^{c},\hat{\textbf{p}}^{s},\hat{\textbf{p}}^{c})}(V)=p_{(g,\textbf{f}^{s},\textbf{f}^{c},\textbf{p}^{s},\textbf{p}^{c})}(V)$ (3)
$\Rightarrow$ $\ \ {\mathbf{z}}_{t}^{s}=(\pi\cdot\mathcal{T})(\hat{g}^{-1}_{s}({\mathbf{x}}_{t})).$
Illustration. Achieving component-wise identification ensures that each estimated style variable corresponds exactly to one true style variable, which facilitates the disentangling of motion features. This enables efficient and precise control over video generation. For example, when generating a video of a person’s face, one can control different aspects, such as head movements or eye blinks, by adjusting the corresponding variables.
@@ 3.1 Block-wise identification
(Linear Operator) Consider two random variables $a$ and $b$ with support $\mathcal{A}$ and $\mathcal{B}$, the linear operator $L_{b\mid a}$ is defined as a mapping from a density function $p_{a}$ in some function space $\mathcal{F}(\mathcal{A})$ onto the density function $L_{b\mid a}\circ p_{a}$ in some function space $\mathcal{F}(\mathcal{B})$,
@@
Illustration of assumptions. The assumptions above are commonly used in the literature on the identification of latent variables under measurement error (hu2008instrumental, ). Firstly, Assumption B1 requires a continuous distribution. Secondly, Assumption B2 imposes a minimal requirement on the number of variables. The linear operator $L_{b|a}$ ensures that there is sufficient variation in the density of $b$ for different values of $a$, thereby guaranteeing injectivity. In a video, ${\mathbf{x}}_{t}$ is of much higher dimensionality compared to the latent variables. As a result, the injectivity assumption is easily satisfied. In practice, following the principle of minimal changes, if a model with fewer latent variables can successfully achieve observational equivalence, it is more likely to learn the true distribution. Assumption B3 requires the distribution of ${\mathbf{x}}_{t}$ changes when the value of latent variables changes. This assumption is much weaker compared to the widely used invertibility assumption adopted by previous works, such as (yao2022temporally, ).
@@
Proof sketch. We separately prove the identifiability of all latent variables and ${\mathbf{z}}^{c}$. For the first part, it is built on fu2025identification , following the line of work from hu2008instrumental . Intuitively, it demonstrates that a minimum of 3 different observations of latent variables are required for identification under the given data generation process. For the second part, we use the contradiction to show that the same ${\mathbf{z}}^{c}$ in different frames of a video can be identified leveraging the invariance.
@@ 3.2 Component-wise identification
Proof sketch. In summary, component-wise identification relies on the changeability of style dynamics, i.e., sufficient changes. Starting from the results of block-wise identification, we establish the connection between ${\mathbf{z}}_{t}^{s}$ and $\hat{{\mathbf{z}}}_{t}^{s}$ in terms of their distributions, i.e., $p({\mathbf{z}}_{t})=p(\hat{{\mathbf{z}}}_{t})\cdot|H_{t}|$, where $H_{t}$ is the jacobian matrix. Leveraging the second-order derivative of the log probability, we construct a system of equations with terms of $\frac{\partial z_{t,i}^{s}}{\partial\hat{z}_{t,j}^{s}}\cdot\frac{\partial z_{t,i}^{s}}{\partial\hat{z}_{t,k}^{s}}$ and coefficients as specified in assumption C2. We leverage the third-order derivative of the previous latent variable $z_{t-1,l}$ to eliminate $|H_{t}|$, utilizing the fact that the history does not influence the current mapping from estimation to truth. Solving this system yields $\frac{\partial z_{t,i}^{s}}{\partial\hat{z}_{t,j}^{s}}\cdot\frac{\partial z_{t,i}^{s}}{\partial\hat{z}_{t,k}^{s}}=0$.
Illustration of assumptions. The assumptions C1 and C2 on the data generating process are commonly adopted in existing identifiable results for Temporally Causal Representation Learning yao2022temporally . Specifically, C1 implies that the latent variables evolve continuously over time, while C2 ensures that the variability in the data can be effectively captured. The assumption C3 constraints to the learned model requires it to separate different variables of style dynamics into independent parts.
@@ 4 Approach
Figure 3: The workflow of the Generator proceeds from left to right, starting with random noise. It first passes through a Temporal Transition Module to obtain a disentangled representation, which is then fed into the synthesis network for frame-level generation. The blue arrow in the figure represents a component-wise flow.
Given our results on identifiability, we implement our model, CoVoGAN. Our architecture is based on StyleGAN2-ADA (Karras2020ada, ), incorporating a Temporal Transition Module in the Generator to enforce the minimal change principle and sufficient change property. Additionally, we add a Video Discriminator to ensure observational equivalence of the joint distribution $p(V)$.
@@ 4.1 Model structure
Temporal Transition Module. Following the generating process in Equation 1, we handle ${\mathbf{z}}^{c}$ and ${\mathbf{z}}_{t}^{s}$ separately. On the one hand, we employ an autoregressive model to capture historical information, followed by a conditional flow to generate ${\mathbf{z}}_{t}^{s}$. Specifically, we implement a Gated Recurrent Unit (GRU) (chung2014empirical, ) and Deep Sigmoid Flow (DSF) (huang2018neural, ), formulated as
${\mathbf{h}}_{t}=\text{GRU}({\mathbf{h}}_{t-1},\epsilon_{t}^{s}),z_{t,i}^{s}=\text{DSF}_{i}(\epsilon_{t,i}^{s};{\mathbf{h}}_{t-1}).$ (5)
On the other hand, since ${\mathbf{z}}^{c}$ are not required to be mutually independent, we use an MLP to generate ${\mathbf{z}}^{c}$, i.e.,
@@
Loss. In addition to the original loss function of StyleGAN2-ADA, we introduce two additional losses: (1) a video discriminator loss, and (2) a mutual information maximization term chen2016infogan between the latent dynamic variables ${\mathbf{z}}_{t}^{s}$ and the intermediate layer outputs of the video discriminator. This encourages the model to learn a more informative and structured representation.
@@
In this subsection, we discuss the relationship between the CoVoGAN  model and the Theorem.
Block-wise identification. As discussed in Theorem 3.2, achieving block-wise identifiability benefits from minimizing the dimension $n_{s}$ of the style dynamics, especially when the true $n_{s}$ is unknown. In practice, this translates to a hyperparameter selection question. In our experiments, we opted for a relatively modest value of $n_{s}$ and observed that it suffices to attain a satisfactory level of disentanglement capability. Furthermore, as required by the assumptions, the learned variables $\hat{{\mathbf{z}}}_{t}^{s}$ and $\hat{{\mathbf{z}}}^{c}$ are block-wise independent, i.e., $\hat{{\mathbf{z}}}_{t}^{s}\perp\!\!\!\perp\hat{{\mathbf{z}}}^{c}$. This independence is necessary to achieve block-wise identifiability.
Component-wise identification. As outlined in Theorem 3.3, the sufficient change property is a critical assumption for achieving identifiability. To enforce temporally conditional independence, we employ a component-wise flow model that transforms a set of independent noise variables $\epsilon_{t}^{s}$ into the style dynamics, conditioned on historical information. Furthermore, the flow model is designed to maximally preserve the information from $\epsilon_{t,i}^{s}$ to $z_{t,i}^{s}$, enabling the model to effectively capture sufficient variability in the data.
Note that when computing the historical information $h_{t}$, we utilize $\epsilon_{t}^{s}$ instead of ${\mathbf{z}}_{t}^{s}$ (as illustrated in the generating process in Equation 1) as the condition for the component-wise flow. This approach offers two key advantages. First, it simplifies the model architecture since the flow does not need to incorporate the output from another flow. Second, the noise terms already fully characterize the corresponding style dynamics, which remains consistent with the theoretical framework.
Furthermore, given that the precise time lag of dynamic variables remains unspecified a priori in the dataset, the GRU’s gating mechanism can selectively filter out irrelevant historical information that lies outside $\text{Pa}({\mathbf{z}}_{t})$. This capability enables the model to demonstrate significantly superior performance compared to traditional non-gated architectures, such as vanilla RNNs.
A detailed ablation study is presented in Section 5.4.

### v3

@@ 1 Introduction
Video generation (Vondrick et al., 2016; Tulyakov et al., 2018; Wang et al., 2022) has become a prominent research focus, driven by its wide-ranging applications in fields such as world simulators (OpenAI, 2024), autonomous driving (Wen et al., 2024; Wang et al., 2023), and medical imaging (Li et al., 2024a; Cao et al., 2024). In particular, controllable video generation (Zhang et al., 2025) is essential for advancing more reliable and efficient video generation models. Despite the impressive results achieved by recent commercial or large-scale models such as Kling (Kuaishou Technology, 2024) and Wan (Wan et al., 2024), precise control over specific aspects of generated video remains a significant challenge, as illustrated in Figure 1. This issue may arise because these models typically represent the video as a unified 4D spatiotemporal block and apply conditioning signals (e.g., text prompts) directly to this global representation. Recent approaches (Ho et al., 2022; Zhou et al., 2022; Yang et al., 2024b; Zheng et al., 2024) follow a similar strategy, with differences in modeling frameworks (e.g., diffusion or VAE) and shapes (e.g., vectors or spatiotemporal blocks). However, these formulations often neglect the intricate spatio-temporal structure of videos, thereby limiting the ability to disentangle and control fine-grained factors, such as head movements and eye blinking.
@@
To address this issue, one intuitive solution is to learn a disentangled representation of the video, within which the internal relationships are often not considered. Some models (Hyvärinen & Oja, 2000; Tulyakov et al., 2018; Yu et al., 2022; Skorokhodov et al., 2022; Wei et al., 2024) explicitly decompose video generation into two parts: motion and identity, representing dynamic and static information, respectively. This separation allows for more targeted control over each aspect, making it possible to modify the motion independently without affecting the identity. (Zhang et al., 2025; Shen et al., 2023) leverage attention mechanisms to further disentangle different concepts within the video, enhancing the ability to control specific features with greater precision. (Fei et al., 2024; Lin et al., 2023) utilize Large Language Models to find the intricate temporal dynamics within the video and then enrich the scene with reasonable details, enabling a more transparent generative process. These methods are intuitive and effective, yet they lack a solid guarantee of disentanglement, making the control less predictable and potentially leading to unintentional coupling of different aspects.
@@
These limitations of previous approaches motivate us to rethink the paradigm of video generation. Inspired by recent advancements in nonlinear Independent Component Analysis (ICA) (Hyvarinen & Morioka, 2017; Khemakhem et al., 2020; Yao et al., 2022; Hyvarinen & Morioka, 2016) and the successful applications like video understanding (Chen et al., 2024), we propose the Controllable Video Generative Adversarial Network (CoVoGAN) with a Temporal Transition Module plugin. Building upon StyleGAN2-ADA (Karras et al., 2020), we distinguish between two types of factors: the dynamic factors that evolve over time, referred to as style dynamics, and the static factors that remain unchanged, which we call content elements. We also distinguish between different components within the dynamics, allowing for more precise control. By leveraging the minimal change principle, we demonstrate their block-wise identifiability (Von Kügelgen et al., 2021; Li et al., 2024b) and find the conditions under which motion and identity can be disentangled, explaining the effectiveness of the previous line of video generation methods that separate motion and identity. In addition, we employ sufficient change property to disentangle different concepts of motion, such as head movement or eye blinking. Specifically, we introduce a flow (Rezende & Mohamed, 2015) mechanism to ensure that the estimated style dynamics are mutually independent conditioned on the historical information. Furthermore, we prove the component-wise identifiability of the style dynamics and provide a disentanglement guarantee for the motion in the video.
@@
We conduct both quantitative and qualitative experiments on various video generation benchmarks. For quantitative evaluation, we use FVD (Fréchet Video Distance) (Unterthiner et al., 2019) to assess the quality of the generated videos. We also compare SAP (Kumar et al., 2017), modularity (Ridgeway & Mozer, 2018) and MCC to demonstrate the disentanglement capability. For qualitative analysis, we evaluate the degree of disentanglement by manipulating different dimensions of the latent variables and comparing the resulting video outputs. Experimental results demonstrate that our method significantly outperforms other video generation models of a scale similar to CoVoGAN.
@@
  • We propose a Temporal Transition Module to achieve a disentangled representation, which leverages the minimal change principle and sufficient change property.
@@
The remainder of this paper is organized as follows. Section 2 formalizes the video generation process and its identifiability, and discusses how this enables disentangled control over video generation. Section 3 provides theoretical insights into when and how block-wise and component-wise identifiability can be achieved. Section 4 introduces the proposed CoVoGAN model and explains how its design is grounded in the theoretical results. Section 5 presents extensive experiments to comprehensively evaluate CoVoGAN. Finally, Section 6 concludes the paper.
@@ 2.1 Generating process
$\left\{\begin{aligned} &{\mathbf{x}}_{t}=g({\mathbf{z}}_{t}^{s},{\mathbf{z}}^{c}),\\ &z_{t,i}^{s}=f^{s}_{i}\left(\mathbf{Pa}(z_{t,i}^{s}),\epsilon_{t,i}^{s}\right),\\ &z_{j}^{c}={f^{c}_{j}\left(\epsilon_{j}^{c}\right)},\end{aligned}\right.\text{with}\left\{\begin{aligned} \epsilon^{s}_{t,i}\sim p^{s}_{\epsilon_{i}},\\ \epsilon^{c}_{j}\sim p^{c}_{\epsilon_{j}},\end{aligned}\right$. (1)
in which $z_{t,i}^{s},z_{j}^{c}\in\mathbb{R}$ refers to the $i$-th entry of ${\mathbf{z}}_{t}^{s}$ and $j$-th entry of ${\mathbf{z}}^{c}$, respectively. $\mathbf{Pa}(z_{t,i}^{s})$ refers to the time-delayed parents of $z_{t,i}^{s}$. All noise terms $\epsilon_{t,i}$ and $\epsilon_{j}$ are independently sampled from their respective distributions: $p_{\epsilon_{i}}^{s}$ for the $i$-th entry of the style dynamics, and $p_{\epsilon_{j}}^{c}$ for the $j$-th entry of the content elements. The components of ${\mathbf{z}}_{t}^{s}$ are mutually independent, conditioned on all historical variables $\cup_{i=1}^{n_{s}}\mathbf{Pa}(z_{t,i}^{s})$. The non-parametric causal transition $f_{i}$ enables an arbitrarily nonlinear interaction between the noise term $\epsilon_{t,i}$ and the set of parent variables $\mathbf{Pa}(z_{t,i}^{s})$, allowing for flexible modeling of the style dynamics.
@@ 2.2 Identification of the latent causal process
$\ \ p_{(\hat{g},\hat{\textbf{f}}^{s},\hat{\textbf{f}}^{c},\hat{\textbf{p}}^{s},\hat{\textbf{p}}^{c})}(V)=p_{(g,\textbf{f}^{s},\textbf{f}^{c},\textbf{p}^{s},\textbf{p}^{c})}(V)\Rightarrow$ $\ \ [{\mathbf{z}}_{t}^{s},{\mathbf{z}}^{c}]=h([\hat{{\mathbf{z}}}_{t}^{s},\hat{{\mathbf{z}}}^{c}]),{\mathbf{z}}^{c}=h_{c}(\hat{{\mathbf{z}}}^{c}),$ (2)
where $\cdot$] denotes concatenation.
@@
$\ \ p_{(\hat{g},\hat{\textbf{f}}^{s},\hat{\textbf{f}}^{c},\hat{\textbf{p}}^{s},\hat{\textbf{p}}^{c})}(V)=p_{(g,\textbf{f}^{s},\textbf{f}^{c},\textbf{p}^{s},\textbf{p}^{c})}(V)\Rightarrow$ $\ \ {\mathbf{z}}_{t}^{s}=(\pi\cdot\mathcal{T})(\hat{g}^{-1}_{s}({\mathbf{x}}_{t})).$ (3)
2.3 From Identifiability to Controllable Video Generation
Identifiability ensures the uniqueness of the latent representation, meaning that the learned latent variables correspond to the true latent variables up to certain allowable transformations. Moreover, identifiability can be defined at different levels, with higher levels indicating that the estimated variables align more closely with the true underlying factors. Stronger identifiability thus leads to more disentangled representations, enabling more efficient and precise control over video generation.
When block-wise identifiability is achieved, content-related components are effectively disentangled from motion. This allows motion control to be applied independently, enabling manipulation of motion without altering the underlying content of the video. For example, in a sequence where a camera moves forward, one can adjust the camera’s direction without affecting the static scene.
In contrast, when component-wise identifiability is achieved, each learned component corresponds one-to-one with its true underlying factor. This property allows independent manipulation of each generative factor by adjusting the corresponding latent dimension, without undesired interference across factors. Such fine-grained control represents the ultimate goal of disentanglement, providing a principled and theoretically grounded formulation. For instance, in generating a video of a face, one can separately adjust head movements or eye blinks by modifying the associated latent variables.
The key remaining question, therefore, is under what conditions identifiability can be guaranteed, and how such guarantees can be established in practice.
@@ 3.1 Block-wise identification
(Linear Operator(Hu & Schennach, 2008a; Dunford & Schwartz, 1988)) Consider two random variables $a$ and $b$ with support $\mathcal{A}$ and $\mathcal{B}$, the linear operator $L_{b\mid a}$ is defined as a mapping from a density function $p_{a}$ in some function space $\mathcal{F}(\mathcal{A})$ onto the density function $L_{b\mid a}\circ p_{a}$ in some function space $\mathcal{F}(\mathcal{B})$,
@@
To better illustrate Linear Operator, examples of linear operators are provided in the Appendix A.3.
@@
Illustration of assumptions. The assumptions above are commonly used in the literature on the identification of latent variables under measurement error (Hu & Schennach, 2008b). Firstly, Assumption B1 requires a continuous distribution. Secondly, Assumption B2 imposes a minimal requirement on the number of variables. The linear operator $L_{b|a}$ ensures that there is sufficient variation in the density of $b$ for different values of $a$, thereby guaranteeing injectivity. In a video, ${\mathbf{x}}_{t}$ is of much higher dimensionality compared to the latent variables. As a result, the injectivity assumption is easily satisfied. In practice, following the principle of minimal changes, if a model with fewer latent variables can successfully achieve observational equivalence, it is more likely to learn the true distribution. Assumption B3 requires the distribution of ${\mathbf{x}}_{t}$ changes when the value of latent variables changes. This assumption is much weaker compared to the widely used invertibility assumption adopted by previous works, such as (Yao et al., 2022).
@@
Proof sketch. We separately prove the identifiability of all latent variables and ${\mathbf{z}}^{c}$. For the first part, it is built on (Fu et al., 2025), following the line of work from (Hu & Schennach, 2008b). Intuitively, it demonstrates that a minimum of 3 different observations of latent variables are required for identification under the given data generation process. For the second part, we use the contradiction to show that the same ${\mathbf{z}}^{c}$ in different frames of a video can be identified leveraging the invariance.
@@ 3.2 Component-wise identification
Proof sketch. In summary, component-wise identification relies on the changeability of style dynamics, i.e., sufficient changes. Starting from the results of block-wise identification, we establish the connection between ${\mathbf{z}}_{t}^{s}$ and $\hat{{\mathbf{z}}}_{t}^{s}$ in terms of their distributions, i.e., $p({\mathbf{z}}_{t})=p(\hat{{\mathbf{z}}}_{t})\cdot|H_{t}|$, where $H_{t}$ is the jacobian matrix. Leveraging the second-order derivative of the log probability, we construct a system of equations with terms of $\frac{\partial z_{t,i}^{s}}{\partial\hat{z}_{t,j}^{s}}\cdot\frac{\partial z_{t,i}^{s}}{\partial\hat{z}_{t,k}^{s}}$ and coefficients as specified in assumption C2. We leverage the third-order derivative of the previous latent variable $z_{t-1,l}$ to eliminate $|H_{t}|$, utilizing the fact that the history does not influence the current mapping from estimation to truth. Solving this system yields $\frac{\partial z_{t,i}^{s}}{\partial\hat{z}_{t,j}^{s}}\cdot\frac{\partial z_{t,i}^{s}}{\partial\hat{z}_{t,k}^{s}}=0$. This indicates that $z^{s}_{t,i}$ is a function of at most one $\hat{z}^{s}_{t,j}$.
Illustration of assumptions. The assumptions C1, C2 on the data generating process are commonly adopted in existing identifiable results for Temporally Causal Representation Learning (Yao et al., 2022). Specifically, C1 implies that the latent variables evolve continuously over time, while C2 ensures that the variability in the data can be effectively captured. The assumption C3 constraints to the learned model require it to separate different variables of style dynamics into independent parts.
@@ 4 Approach
Figure 3: Generator operates from left to right, beginning with a random noise input. The noise first passes through a Temporal Transition Module, which produces a disentangled representation of the underlying factors. This representation is then fed into the synthesis network to generate frames at the pixel level. In the figure, the blue arrow illustrates the Deep Sigmoid Flow.
Given our results on identifiability, we implement our model, CoVoGAN. Our architecture is based on StyleGAN2-ADA (Karras et al., 2020), incorporating a Temporal Transition Module in the Generator to enforce the minimal change principle and sufficient change property. Additionally, we add a Video Discriminator to ensure observational equivalence of the joint distribution $p(V)$.
@@ 4.1 Model structure
Temporal Transition Module. Following the generating process in Equation 1, we handle ${\mathbf{z}}^{c}$ and ${\mathbf{z}}_{t}^{s}$ separately. On the one hand, we employ an autoregressive model to capture historical information, followed by a conditional flow to generate ${\mathbf{z}}_{t}^{s}$. Specifically, we implement a Gated Recurrent Unit (GRU) (Chung et al., 2014) and Deep Sigmoid Flow (DSF) (Huang et al., 2018), formulated as
${\mathbf{h}}_{t}=\text{GRU}({\mathbf{h}}_{t-1},\epsilon_{t-1}^{s}),z_{t,i}^{s}=\text{DSF}_{i}(\epsilon_{t,i}^{s};{\mathbf{h}}_{t-1}),$ (5)
where ${\mathbf{h}}_{t}$ denotes all the historical information until time step $t$, and $z_{t,i}^{s}$ will be mutually independent conditioned on ${\mathbf{h}}_{t}$. On the other hand, since ${\mathbf{z}}^{c}$ are not required to be mutually independent, we use an MLP to generate ${\mathbf{z}}^{c}$, i.e.,
@@
Loss. In addition to the original loss function of StyleGAN2-ADA, we introduce two additional losses: (1) a video discriminator loss, and (2) a mutual information maximization term (Chen et al., 2016) between the latent dynamic variables ${\mathbf{z}}_{t}^{s}$ and the intermediate layer outputs of the video discriminator. This encourages the model to learn a more informative and structured representation.
@@
Block-wise identification. As discussed in Theorem 3.2, achieving block-wise identifiability benefits from minimizing the dimension $n_{s}$ of the style dynamics, especially when the true $n_{s}$ is unknown. In practice, this translates to a hyperparameter selection question. Therefore, we opted for a relatively modest value of $n_{s}$ and observed that it suffices to attain a satisfactory level of disentanglement capability. Furthermore, as required by the assumptions, the learned variables $\hat{{\mathbf{z}}}_{t}^{s}$ and $\hat{{\mathbf{z}}}^{c}$ are block-wise independent, i.e., $\hat{{\mathbf{z}}}_{t}^{s}\perp\!\!\!\perp\hat{{\mathbf{z}}}^{c}$. This independence is necessary to achieve block-wise identifiability.
Component-wise identification. As outlined in Theorem 3.3, the sufficient change property is a critical assumption for achieving identifiability. To enforce temporally conditional independence, we employ a component-wise flow model that transforms a set of independent noise variables $\epsilon_{t}^{s}$ into the style dynamics, conditioned on historical information. Furthermore, the flow model is designed to maximally preserve the information from $\epsilon_{t,i}^{s}$ to $z_{t,i}^{s}$, enabling the model to effectively capture sufficient variability in the data. Note that when computing the historical information $h_{t}$, we utilize $\epsilon_{t}^{s}$ instead of ${\mathbf{z}}_{t}^{s}$ (as illustrated in the generating process in Equation 1) as the condition for the component-wise flow. This approach offers two key advantages. First, it simplifies the model architecture since the flow does not need to incorporate the output from another flow. Second, the noise terms already fully characterize the corresponding style dynamics, which remains consistent with the theoretical framework. Furthermore, given that the precise time lag of dynamic variables remains unspecified a priori in the dataset, the GRU’s gating mechanism can selectively filter out irrelevant historical information that lies outside $\text{Pa}({\mathbf{z}}_{t})$. This capability enables the model to demonstrate significantly superior performance compared to traditional non-gated architectures, such as vanilla RNNs. A detailed ablation study is presented in Section 5.4.